OpenStudy (anonymous):

What is the simplified form of 4z^2 - 4z - 15 over 2Z^2 + z - 15

OpenStudy (anonymous):

4z^2 - 4z - 15 =4z^2-10z+6z-15 =2z(2z-5)+3(2z-5) =(2z-5)(2z+3)

OpenStudy (anonymous):

2Z^2 + z - 15 =2z^2+6z-5z-15 =2z(z+3)-5(z+3) =(z+3)(2z-5)